what line passes through 8,10 and is parallel to the graph of y=8/3x+7

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

y = 2 2/3 x + 7
Compare this equation with y=mx+b
slope m = 8/3
The slope of a line parallel to the above line will be the same
The slope of the required line will be 8/3
m= 8/3 ,point ( 8 , 10 )
Find b by plugging the values of m & the point in
y=mx+b
10 = 64/ 3 + b
b= -34/3
m= 8/3
Plug value of the slope and b in y = mx +b
The required equation is y = 8/3 x -34/ 3

m.ananth@hotmail.ca

Answer by stanbon(75887)

what line passes through 8,10 and is parallel to the graph of y=8/3x+7
-----
slope of the line must be (8/3).
--------------
Form: y = mx + b
You know y = 10 when x = 8, and m = 8/3 ; solve for "b":
---------
10 = (3/8)8 + b
b = 10-3
b = 7
-----
Equation:
y = (8/3)x+7
=======================
Cheers,
Stan H.